Sales at LL Boutique decreased 10% this month compared to last month. If sales this month were $139,581, what were the sales (in $) last month?



1
Expert's answer
2021-02-24T07:24:21-0500

Sales figure for last month = "x"

Sales figure for this month = $139581

Since sales dropped by 10% compared to last month, then

10% = 0.1


"\frac{x - 139581{}}{x} = 0.1"

"x{-139581{}}{} = 0.1x"

"{x-0.1x{}} = 139581"

"{0.9x{}} = 139581"

= $155090

The sales figure for last month were $155090


"{check{}}:"

"\frac{155090 -139581{}}{155090} ×100"


"=\frac{15509{}}{155090}×100"

= 10%